OpenStudy (anonymous):

: If a body travels 0.2m in the first 0.2s and 0.22m in the next 0.4s , what will be its velocity at the end of the 0.7th second ?

OpenStudy (amistre64):

are we to assume a constant acceleration?

OpenStudy (amistre64):

given to points to work from, the simplest thing you can get is a linear function

OpenStudy (amistre64):

( t, m) ( .2, .2) -(.4,.22) ------- -.2, -.02; slope between them is 1/10

OpenStudy (amistre64):

y = .1t -.1(.2) +.2 y = .1t -.02 +.2 y = .1t + .18 when t = .7 we get: y = .1(.7) + .18 = .25 maybe

OpenStudy (amistre64):

these are positions tho, and not velocity. the velocity itself is just .1m/s i believe

OpenStudy (amistre64):

but im reading it wrong again i believe ..... If a body travels 0.2m in the first 0.2s and so at .2 we have a position of .2m 0.22m in the next 0.4s at .6s we have a position of .22m or an extra .22m?

OpenStudy (amistre64):

still, given 2 data points, the best we can assume is a linear model and the velocity at any given point is the same as the slope of the line

OpenStudy (amistre64):

but then again we have an extra point dont we, i assume at time 0s we are at a velocity of 0

OpenStudy (amistre64):

if we are measureing it from rest then we need to fit a curve to 3 points: (0,0), (.2,.2), and (.6,.42) perhaps

OpenStudy (amistre64):

our curve fits are then, if im reading the information correctly: a0^2 + b0 + c = 0 a.2^2 + b.2 + c = .2 a..6^2 + b.6 + c = .42 c = 0 .04a + .2b = .2 .36a + .6b = .42 this turns into a doable system of equations by applying cramers rule: b = .42(.04)-.36(.2) / .6(.04)-.36(.2) = 1.15 a = .42(.2)-.6(.2) / .36(.2)- .6(.04) = -.75 so our curve should be: -.75 t^2 + 1.15 t ; let me dbl chk this with the wolf tho ... the wolf like it

OpenStudy (amistre64):

with this as the position equation, we can take the derivative to determine the instanteous velocity at any given point. v(t) = -1.5 t + 1.15 v(.7) = .1 m/s

OpenStudy (anonymous):

i think you can work this out in terms of the average velocity of the body in each interval (assumong constant acceleration) we know s=d/t so therefore average speed for first 0.2 seconds= 1m/s therefore its maximum speed must be 2m/s (at 0.2s, at 0s v=o). now fro the second interval, average speed = 2.2/0.2=11 m/s since you started at 1m/s and get an average of 11 velocity at 0.4s =21m/s. therefore the acceleration is (21-2)/0.2=95m/s^2. therefore at 0.7s v= 95x0.7=66.5m/s. this seems right to me but the previous answer is just as likely to be correct if not more so considering i'm the one giving my solution :p.